OpenStudy (anonymous):

how do you find the range of f(x) = (x + 5)^2 + 8

OpenStudy (trojanpoem):

I assume, you're looking for an algebraic method , as it's easy in graphs. Find the inverse of this function: Let y = f(x) y = (x+5)^2 + 8 y - 8 = (x+5)^2 sqrt(y - 8) - 5 = x x = sqrt(y - 8) - 5 y-8 >= 0 y >= 8 ]inf, 8]

OpenStudy (phi):

the range are all the f(x) values that you can get as you change x it helps to know the range of f(x)= x^2 do you know that x^2 is 0 or positive, not matter what x is ?

OpenStudy (anonymous):

@TrojanPoem that helps a lot. Thank you.

OpenStudy (phi):

similarly , the smallest (x+5)^2 can be is 0 so the smallest f(x) can be is 0+8= 8 and of course, you can get as big a number as you like if you make x big enough so [8, +infinity)

OpenStudy (trojanpoem):

Any time.